# Catching the Details: Self-Distilled RoI Predictors for Fine-Grained MLLM Perception

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) require high-resolution visual information to perform fine-grained perception, yet processing entire high-resolution images is computationally prohibitive. While recent methods leverage a Region-of-Interest (RoI) mechanism to focus on salient areas, they typically present a difficult trade-off: training-based approaches depend on large-scale annotated datasets, while training-free methods that utilize the model’s internal attention are computationally inefficient and less accurate, requiring either multi-pass prefill stages or reliance on the slow auto-regressive decoding process. In this paper, we propose an efficient, annotation-free Self-Distilled Region Proposal Network (SD-RPN) that resolves this trade-off. The SD-RPN is built around a pipeline that transforms the noisy attention maps from the MLLM’s middle layers into high-quality pseudo-RoI labels by explicitly denoising the signal and resolving ambiguity. We use these labels to train a lightweight Region Proposal Network (RPN) that learns a more precise localization. This RPN is also highly efficient, predicting the RoI in a single forward pass using features from the MLLM’s middle layers, decoupling RoI identification from the auto-regressive generation and avoiding costly multi-pass operations. To validate our approach, we integrate the framework into multiple MLLM families. Despite being trained on only a few (e.g. 10K) question-answer pairs, our method demonstrates exceptional data efficiency and generalization, achieving over a 10% absolute accuracy improvement on unseen benchmarks, including TextVQA, DocVQA, and V-Star. Our work presents a practical and scalable solution for enhancing the fine-grained perception of MLLMs without requiring costly supervision or full model fine-tuning. Code is available at https://github.com/YuHengsss/SD-RPN.
@@ 1 Introduction
Recent years have witnessed significant advancements in Multimodal Large Language Models (MLLMs), which have evolved from foundational architectures like LLaVA (Liu et al., 2023b) to more sophisticated systems (Wang et al., 2024a; Bai et al., 2023; Chen et al., 2024a; b) such as Qwen2.5-VL (Bai et al., 2025) and InternVL-3.0 (Zhu et al., ). The performance of MLLMs is tied to the quality of their visual perception. In a typical MLLM architecture, a vision encoder, such as CLIP (Radford et al., 2021), processes visual signals and projects them into the embedding space of a Large Language Model (LLM) (Touvron et al., 2023; Zheng et al., 2023) for subsequent reasoning. Consequently, the richness of these visual features is crucial for the model to achieve a comprehensive and fine-grained understanding of the input.
@@
To enhance this fine-grained perception, scaling the resolution of visual inputs has emerged as a direct and effective strategy. The approaches to achieve this have evolved considerably. Initial methods (Liu et al., 2023b; Dai et al., 2023; Bai et al., 2023) relied on a fixed-resolution input. Subsequent research introduced more flexible techniques, such as S2 (Shi et al., 2024), Any-Resolution (Liu et al., 2023a; Chen et al., 2024b), and Naive Dynamic Resolution (Wang et al., 2024a; Bai et al., 2025) to handle higher-resolution imagery. However, processing entire high-resolution images uniformly is computationally intensive. More recently, an alternative paradigm (Yu et al., 2025; Shi et al., 2025) has gained traction: identifying a Region-of-Interest (RoI) within a low-resolution input and then selectively scaling up only that specific region. This RoI-centric approach has proven to be a more efficient and effective means of improving visual detail perception.
@@
Figure 1: (a) The Pipeline of SD-RPN. The RPN is trained with pseudo-labels to effectively predict RoIs. These RoIs are then used to crop fine-grained sub-images for the final inference stage. (b) Performance Comparison. Performance evaluation with S2 (Shi et al., 2024) and ViCrop (Zhang et al., 2025) on the LLaVA-1.5-7B baseline. Accuracy is averaged over five Document and OCR benchmarks. Our SD-RPN achieves a superior trade-off between performance and throughput.
@@
While the RoI paradigm marks a significant step toward efficient high-resolution processing, current methodologies for identifying these regions still present notable limitations. Approaches such as VILA-HD (Shi et al., 2025), for instance, rely on large-scale pre-training with detailed annotations, a process that is both data-intensive and computationally demanding. Furthermore, they often require a complete prefilling stage for the initial low-resolution image, which can impede both training and inference efficiency. An alternative direction leverages the intrinsic localization capabilities of MLLMs (Kang et al., 2025b; Zhang et al., 2025) without training, specifically by computing cross-attention scores between image tokens and the corresponding textual description. However, leveraging this internal attention for RoI identification is often computationally inefficient in inference, as current methods typically require either complex, multi-pass operations during the prefill stage (Zhang et al., 2025) or rely on the inherently sequential and slow auto-regressive decoding stage (Wang et al., 2024b). Consequently, an effective and efficient method for RoI identification that avoids both reliance on extensive annotated data and the high latency of auto-regressive decoding or multiple forward passes remains a critical, yet underexplored, challenge.
@@
In this work, we introduce SD-RPN, a novel self-distillation framework (as shown in Fig. 1(a)) designed to overcome these limitations by efficiently harnessing the MLLM’s intrinsic localization capabilities. Our approach is motivated by the insight that while an MLLM’s internal attention provides a strong RoI signal, it is too noisy (e.g., attention sinks, incomplete activation) for direct supervision. Previous studies reveal that using such noisy signals for dense supervision yields sub-optimal results (Wang et al., 2022; Tian et al., 2021). To address this, we propose a pseudo-labeling pipeline that transforms the noisy attention map into a sparse and reliable supervisory signal. This pipeline first denoises the map by removing sink tokens based on feature norms. Subsequently, it employs a selective classification strategy that assigns discrete labels based on confidence thresholds and a minimal bounding box around high-attention tokens, which resolves the ambiguity inherent in the original noisy map. We use these pseudo labels to train a lightweight RPN. By learning from this distilled knowledge, the RPN develops a more precise localization function than methods relying on raw, ambiguous attention. Beyond its accuracy, the RPN’s architectural design confers significant efficiency gains. Composed of a few transformer blocks built upon the frozen MLLM backbone, the RPN operates on features from the model’s middle layers. This strategic placement allows it to predict the RoI by executing only a partial forward pass up to the MLLM’s middle layers, completely decoupling localization from the slow, auto-regressive generation process. The entire framework is trained end-to-end, distilling the localization knowledge from the model’s own response-guided attention into the efficient RPN. This enables a dynamic two-stage inference process where the model first predicts salient regions and then analyzes high-resolution crops to generate its final response.
@@
Our SD-RPN achieves a significant enhancement (Fig. 1(b)) in fine-grained perception without the cost of full-model fine-tuning or large-scale annotated datasets. Moreover, our framework demonstrates exceptional training efficiency and generalization. To validate our method, we integrate it into the widely-used LLaVA-1.5 (Liu et al., 2023a), aligning with previous works (Zhang et al., 2025; Kang et al., 2025b). When being trained on only 10K randomly selected question-answer pairs from GQA (Hudson & Manning, 2019) and OCR-VQA (Mishra et al., 2019), our method yields substantial gains on unseen data, achieving over a 10% absolute accuracy improvement on benchmarks such as TextVQA (Singh et al., 2019), DocVQA (Mathew et al., 2020), and V-Star (Wu & Xie, 2023). To further demonstrate the broad applicability of our approach, we also extend our validation to the more recent DeepSeek-VL (Lu et al., 2024) and Qwen2.5-VL (Bai et al., 2025) architectures. Our contributions are twofold. First, we introduce a robust pipeline to denoise the internal attention maps of an MLLM, generating high-quality pseudo-labels for supervision. Second, we propose a novel, annotation-free self-distillation framework that trains a lightweight RPN to predict RoIs by leveraging the MLLM’s intrinsic localization knowledge.
@@ 2 Related Works
Recent studies have established that MLLMs often struggle with fine-grained perception, a limitation rooted in the challenge of efficiently processing high-resolution visual inputs (Tong et al., 2024b; Kang et al., 2025b; Shi et al., 2024; Liu et al., 2024a; Chen et al., 2024a). One line of research has focused on enhancing the model’s global visual understanding. This includes developing more sophisticated vision encoders (Wang et al., 2024a; Bai et al., 2025; Zhang et al., 2024b; Ge et al., 2024; Luo et al., 2024), supplementing the LLM with full high-resolution images (Liu et al., 2024a; Zhu et al., ; Huang et al., 2024; Tong et al., 2024a), and incorporating external tools (Zhao et al., 2024). More recent works Shi et al. (2025); Yu et al. (2025); Zheng et al. (2025); Shao et al. (2024) have demonstrated that identifying the RoI first in relatively low-resolution visual input and then scaling the resolution specifically for the RoI is more efficient and effective. However, they come at the expense of extensive training, requiring massive supervision and costly annotations. While recent training-free methods (Zhang et al., 2025; Wang et al., 2024b) attempt to identify RoIs by utilizing the internal perceptual capabilities of MLLMs, they are often hampered by noisy activations or require slow auto-regresssive decoding stage and multiple forward passes, which hinders both performance and efficiency.
@@ 3.2 Pseudo-Label Generation for RoI
The first source of noise we address is the sink tokens. These are visual tokens that attract substantial attention despite being semantically irrelevant to the grounded object. As identified in recent studies (Darcet et al., 2023; Kang et al., 2025a), sink tokens can be identified by the high L2-norm of their corresponding feature vectors $\mathbf{H}_{v}$. To mitigate them, we first denoise the initial RoI map $\mathbf{M}_{\text{RoI}}$ by nullifying the scores of all identified sink tokens, yielding a cleaner version, $\mathbf{M}^{\prime}_{\text{RoI}}$:
@@
where $\tau_{\text{norm}}$ is a predefined norm threshold.
@@
Following the removal of sink tokens, the denoised RoI map $\mathbf{M}^{\prime}_{\text{RoI}}$ more clearly highlights the foreground area. However, it is not yet an ideal supervisory signal, as it still suffers from an obscure foreground-background margin and incomplete object activation.
To empirically investigate the margin issue, we analyze the denoised attention maps $\mathbf{M}^{\prime}_{\text{RoI}}$ on the TextVQA subset (Zhang et al., 2025). Fig. 3 plots the proportion of tokens falling inside the ground-truth (GT) bounding box as a function of their sample-wise relative attention score, calculated as $a/a_{\max}$ for each attention value $a\in\mathbf{M}^{\prime}_{\text{RoI}}$. The histogram reveals that the proportion of tokens inside the GT box is approximately 40% for high attention scores (e.g., >0.2>0.2) and 10% for low scores (e.g., <0.1<0.1). However, the ambiguous middle-range of attention not only exhibit a distinct localization pattern from the more decisive high and low scores, but they also vastly outnumber the high-attention tokens. Using such a signal for dense supervision would inevitably force the model to learn from these numerous ambiguous regions. To address this, we avoid regressing on $\mathbf{M}^{\prime}_{\text{RoI}}$ directly and instead implement a selective binary classification that assigns foreground or background only to high-confidence tokens, leaving ambiguous tokens to be ignored. This approach is also designed to alleviate the incomplete activation problem. We define a minimal bounding box, $\mathcal{B}_{fg}$, which encloses identified foreground tokens. Tokens inside this box that is not classified as foreground is explicitly ignored, which prevents the RPN from receiving erroneous background signals from inactivated parts of the true object. Guided by these criteria, the final pseudo-label map, $\bar{\mathbf{M}}_{\text{RoI}}$ is constructed by:
@@
where the foreground set is $\mathcal{S}_{fg}=\{j\mid a_{j}\geq\tau_{fg}\,a_{\max}\}$ and the background set is $\mathcal{S}_{bg}=\{j\mid j\notin\mathcal{B}_{fg}\text{ and }a_{j}\leq\tau_{bg}\,a_{\max}\}$. In Appendix E, we also provide a theoretical view of why learning to predict RoI labels is superior than using the raw attention maps directly.
@@ 3.3 RoI Prediction via Self-distillation
where Norm denotes the layer normalization (Zhang & Sennrich, 2019; Ba et al., 2016). The predicted dense RoI map, $\hat{\mathbf{M}}_{\text{RoI}}$ is then computed via a simple matrix multiplication of these query and key matrices: $\hat{\mathbf{M}}_{\text{RoI}}=\mathbf{Q}_{\text{RoI}}\mathbf{K}_{v}^{T}$. For brevity, the head dimension is omitted here; in practice, attention scores are computed per head and then averaged. Notably, this design is more efficient than exsiting frameworks (Shi et al., 2025; Zhang et al., 2025) as it only need a single forward through partial LLM layers.
@@
The RPN is trained entirely through self-distillation to predict the pseudo-label map $\bar{\mathbf{M}}_{\text{RoI}}$ which is generated via the pipeline in Section 3.2 by minimizing a binary cross-entropy (BCE) loss, defined as $\mathcal{L}_{\text{BCE}}(\hat{\mathbf{M}}_{\text{RoI}},\bar{\mathbf{M}}_{\text{RoI}})$. While the text responses used for pseudo-label generation could theoretically originate from a stronger teacher model (e.g., GPT-4) or human annotations, our empirical results reveal a crucial insight: using the MLLM’s self-predicted responses yields superior performance (detailed in Tab. 4(b)). We attribute this counter-intuitive result to the principle of representational consistency. Although an external teacher’s response may be more accurate, the attention maps it induces can be out-of-distribution for the student RPN. In contrast, the attention maps from self-generated responses, even if imperfect, are inherently aligned with its own internal visual grounding mechanisms. This creates a more consistent and attainable distillation target, which proves more effective for training the lightweight RPN, especially in our data-efficient setting. This self-sufficient framework thus removes any dependency on external models or annotated data.

### v3

@@ Abstract
Multimodal Large Language Models (MLLMs) require high-resolution visual information to perform fine-grained perception, yet processing entire high-resolution images is computationally prohibitive. While recent methods leverage a Region-of-Interest (RoI) mechanism to focus on salient areas, they typically present a difficult trade-off: training-based approaches depend on large-scale annotated datasets, while training-free methods that utilize the model’s internal attention are computationally inefficient and less accurate, requiring either multi-pass prefill stages or reliance on the slow auto-regressive decoding process. In this paper, we propose an efficient, annotation-free Self-Distilled Region Proposal Network (SD-RPN) that resolves this trade-off. The SD-RPN is built around a pipeline that transforms the noisy attention maps from the MLLM’s middle layers into high-quality pseudo-RoI labels by explicitly denoising the signal and resolving ambiguity. We use these labels to train a Region Proposal Network (RPN) that learns a more precise localization. This RPN is also highly efficient, predicting the RoI in a single forward pass using features from the MLLM’s middle layers, decoupling RoI identification from the auto-regressive generation and avoiding costly multi-pass operations. To validate our approach, we integrate the framework into the LLaVA-1.5 architecture. Despite being trained on only a few (e.g. 10K) question-answer pairs, our method demonstrates exceptional data efficiency and generalization, achieving over a 10% absolute accuracy improvement on unseen benchmarks, including TextVQA, DocVQA, and V-Star. Our work presents a practical and scalable solution for enhancing the fine-grained perception of MLLMs without requiring costly supervision or full model fine-tuning. Code is available at https://github.com/YuHengsss/SD-RPN.
@@ 1 Introduction
Recent years have witnessed significant advancements in Multimodal Large Language Models (MLLMs), which have evolved from foundational architectures like LLaVA (Liu et al., 2023b) to more sophisticated systems (Wang et al., 2024a; Bai et al., 2023; Chen et al., 2024a; b) such as Qwen2.5-VL (Bai et al., 2025) and InternVL-3.0 (Zhu et al., 2025). The performance of MLLMs is tied to the quality of their visual perception. In a typical MLLM architecture, a vision encoder, such as CLIP (Radford et al., 2021), processes visual signals and projects them into the embedding space of a Large Language Model (LLM) (Touvron et al., 2023; Zheng et al., 2023) for subsequent reasoning. Consequently, the richness of these visual features is crucial for the model to achieve a comprehensive and fine-grained understanding of the input.
@@
To enhance this fine-grained perception, scaling the resolution of visual inputs has emerged as a direct and effective strategy. The approaches to achieve this have evolved considerably. Initial methods (Liu et al., 2023b; Dai et al., 2023; Bai et al., 2023) relied on a fixed-resolution input. Subsequent research introduced more flexible techniques, such as S2 (Shi et al., 2024), Any-Resolution (Liu et al., 2023a; Chen et al., 2024b), and Naive Dynamic Resolution (Wang et al., 2024a; Bai et al., 2025) to handle higher-resolution imagery. However, processing entire high-resolution images uniformly is computationally intensive. More recently, an alternative paradigm (Yu et al., 2025; Shi et al., 2025a) has gained traction: identifying a Region-of-Interest (RoI) within a low-resolution input and then selectively scaling up only that specific region. This RoI-centric approach has proven to be a more efficient and effective means of improving visual detail perception.
@@
Figure 1: (a) The Pipeline of SD-RPN. The RPN is trained with pseudo-labels to effectively predict RoIs. These RoIs are then used to crop fine-grained sub-images for the final inference stage. (b) Performance Comparison. Performance evaluation with S2 (Shi et al., 2024) and ViCrop (Zhang et al., 2025a) on the LLaVA-1.5-7B baseline. Accuracy is averaged over five Document and OCR benchmarks. Our SD-RPN achieves a superior trade-off between performance and throughput.
@@
While the RoI paradigm marks a significant step toward efficient high-resolution processing, current methodologies for identifying these regions still present notable limitations. Approaches such as VILA-HD (Shi et al., 2025a), for instance, rely on large-scale pre-training with detailed annotations, a process that is both data-intensive and computationally demanding. Furthermore, they often require a complete prefilling stage for the initial low-resolution image, which can impede both training and inference efficiency. An alternative direction leverages the intrinsic localization capabilities of MLLMs (Kang et al., 2025b; Zhang et al., 2025a) without training, specifically by computing cross-attention scores between image tokens and the corresponding textual description. However, leveraging this internal attention for RoI identification is often computationally inefficient in inference, as current methods typically require either complex, multi-pass operations during the prefill stage (Zhang et al., 2025a) or rely on the inherently sequential and slow auto-regressive decoding stage (Wang et al., 2024b). Consequently, an effective and efficient method for RoI identification that avoids both reliance on extensive annotated data and the high latency of auto-regressive decoding or multiple forward passes remains a critical, yet underexplored, challenge.
@@
In this work, we introduce SD-RPN, a novel self-distillation framework (as shown in Fig. 1(a)) designed to overcome these limitations by efficiently harnessing the MLLM’s intrinsic localization capabilities. Our approach is motivated by the insight that while an MLLM’s internal attention provides a strong RoI signal, it is too noisy (e.g., attention sinks, incomplete activation) for direct supervision. Previous studies reveal that using such noisy signals for dense supervision yields sub-optimal results (Wang et al., 2022; Tian et al., 2021). To address this, we propose a pseudo-labeling pipeline that transforms the noisy attention map into a sparse and reliable supervisory signal. This pipeline first denoises the map by removing sink tokens based on feature norms. Subsequently, it employs a selective classification strategy that assigns discrete labels based on confidence thresholds and a minimal bounding box around high-attention tokens, which resolves the ambiguity inherent in the original noisy map. We use these pseudo labels to train a RPN. By learning from this distilled knowledge, the RPN develops a more precise localization function than methods relying on raw, ambiguous attention. Beyond its accuracy, the RPN’s architectural design confers significant efficiency gains. Composed of a few transformer blocks built upon the frozen MLLM backbone, the RPN operates on features from the model’s middle layers. This strategic placement allows it to predict the RoI by executing only a partial forward pass up to the MLLM’s middle layers, completely decoupling localization from the slow, auto-regressive generation process. The entire framework is trained end-to-end, distilling the localization knowledge from the model’s own response-guided attention into the efficient RPN. This enables a dynamic two-stage inference process where the model first predicts salient regions and then analyzes high-resolution crops to generate its final response.
@@
Our SD-RPN achieves a significant enhancement (Fig. 1(b)) in fine-grained perception without the cost of full-model fine-tuning or large-scale annotated datasets. Moreover, our framework demonstrates exceptional training efficiency and generalization. Using only 10K random samples from GQA (Hudson and Manning, 2019) and OCR-VQA (Mishra et al., 2019) within the LLaVA-1.5 (Liu et al., 2023a) framework, our approach yields substantial gains on unseen data, boosting accuracy by over 10% on TextVQA (Singh et al., 2019), DocVQA (Mathew et al., 2020), and V-Star (Wu and Xie, 2023). We further confirm the robustness of our method by applying it to the more recent DeepSeek-VL (Lu et al., 2024) and Qwen2.5-VL (Bai et al., 2025) architectures. Our contributions are twofold. First, we introduce a robust pipeline to denoise the internal attention maps of an MLLM, generating high-quality pseudo-labels for supervision. Second, we propose a novel, annotation-free self-distillation framework that trains a RPN to predict RoIs by leveraging the MLLM’s intrinsic localization knowledge.
@@ 2 Related Works
Recent studies have established that MLLMs often struggle with fine-grained perception, a limitation rooted in the challenge of efficiently processing high-resolution visual inputs (Tong et al., 2024b; Kang et al., 2025b; Shi et al., 2024; Liu et al., 2024a; Chen et al., 2024a). One line of research has focused on enhancing the model’s global visual understanding. This includes developing more sophisticated vision encoders (Wang et al., 2024a; Bai et al., 2025; Zhang et al., 2024b; Ge et al., 2024; Luo et al., 2024), supplementing the LLM with full high-resolution images (Liu et al., 2024a; Zhu et al., 2025; Huang et al., 2024; Tong et al., 2024a; Cha et al., 2024; Li et al., 2025; 2024b), and incorporating external tools (Zhao et al., 2024). More recent works Shi et al. (2025a); Yu et al. (2025); Zheng et al. (2025); Shao et al. (2024) have demonstrated that identifying the RoI first in relatively low-resolution visual input and then scaling the resolution specifically for the RoI is more efficient and effective. However, they come at the expense of extensive training, requiring massive supervision and costly annotations. While recent training-free methods (Zhang et al., 2025a; Wang et al., 2024b; Shi et al., 2025b) attempt to identify RoIs by utilizing the internal perceptual capabilities of VLMs, they are often hampered by noisy activations or require slow auto-regresssive decoding stage and multiple forward passes, which hinders both performance and efficiency.
@@ 3.2 Pseudo-Label Generation for RoI
The first source of noise we address is the sink tokens. These are visual tokens that attract substantial attention despite being semantically irrelevant to the grounded object. As identified in recent studies (Darcet et al., 2023; Kang et al., 2025a), sink tokens can be identified by the high L2-norm of their corresponding feature vectors $\mathbf{H}_{v}$. To mitigate them, we first denoise the initial RoI map $\mathbf{M}_{\text{RoI}}$ suppressing sink tokens identified via a predefined norm threshold $\tau_{\text{norm}}$, yielding a cleaner version, $\mathbf{M}^{\prime}_{\text{RoI}}$:
@@
Following the removal of sink tokens, the denoised RoI map $\mathbf{M}^{\prime}_{\text{RoI}}$ more clearly highlights the foreground area. However, it is not yet an ideal supervisory signal, as it still suffers from an obscure foreground-background margin and incomplete object activation. To empirically investigate the margin issue, we analyze the denoised attention maps $\mathbf{M}^{\prime}_{\text{RoI}}$ on the TextVQA subset (Zhang et al., 2025a). Fig. 3 plots the proportion of tokens falling inside the ground-truth (GT) bounding box as a function of their sample-wise relative attention score, calculated as $a/a_{\max}$ for each attention value $a\in\mathbf{M}^{\prime}_{\text{RoI}}$. The histogram reveals that the proportion of tokens inside the GT box is approximately 40% for high attention scores (e.g., >0.2>0.2) and 10% for low scores (e.g., <0.1<0.1). However, the ambiguous middle-range of attention not only exhibit a distinct localization pattern from the more decisive high and low scores, but they also vastly outnumber the high-attention tokens. Using such a signal for dense supervision would inevitably force the model to learn from these numerous ambiguous regions. To address this, we avoid regressing on $\mathbf{M}^{\prime}_{\text{RoI}}$ directly and instead implement a selective binary classification that assigns foreground or background only to high-confidence tokens, leaving ambiguous tokens to be ignored. This approach is also designed to alleviate the incomplete activation problem. We define a minimal bounding box, $\mathcal{B}_{fg}$, which encloses identified foreground tokens. Tokens inside this box that is not classified as foreground is explicitly ignored, which prevents the RPN from receiving erroneous background signals from inactivated parts of the true object. Guided by these criteria, the final pseudo-label map, $\bar{\mathbf{M}}_{\text{RoI}}$ is constructed by:
@@
where the foreground set is $\mathcal{S}_{fg}=\{j\mid a_{j}\geq\tau_{fg}\,a_{\max}\}$ and the background set is $\mathcal{S}_{bg}=\{j\mid j\notin\mathcal{B}_{fg}\text{ and }a_{j}\leq\tau_{bg}\,a_{\max}\}$. In Appendix F, we also provide a theoretical view of why learning to predict RoI labels is superior than using the raw attention maps directly.
@@ 3.3 RoI Prediction via Self-distillation
where Norm denotes the layer normalization (Zhang and Sennrich, 2019; Ba et al., 2016). The predicted dense RoI map, $\hat{\mathbf{M}}_{\text{RoI}}$ is then computed via a simple matrix multiplication of these query and key matrices: $\hat{\mathbf{M}}_{\text{RoI}}=\mathbf{Q}_{\text{RoI}}\mathbf{K}_{v}^{T}$. For brevity, the head dimension is omitted here; in practice, attention scores are computed per head and then averaged. Notably, this design is more efficient than exsiting frameworks (Shi et al., 2025a; Zhang et al., 2025a) as it only need a single forward through partial LLM layers.
@@
The RPN is trained entirely through self-distillation to predict the pseudo-label map $\bar{\mathbf{M}}_{\text{RoI}}$ which is generated via the pipeline in Section 3.2 by minimizing a binary cross-entropy (BCE) loss, defined as $\mathcal{L}_{\text{BCE}}(\hat{\mathbf{M}}_{\text{RoI}},\bar{\mathbf{M}}_{\text{RoI}})$. While the text responses used for pseudo-label generation could theoretically originate from a stronger teacher model (e.g., GPT-4) or human annotations, our empirical results reveal a crucial insight: using the MLLM’s self-predicted responses yields superior performance (detailed in Tab. 4(b)). We attribute this counter-intuitive result to the principle of representational consistency. Although an external teacher’s response may be more accurate, the attention maps it induces can be out-of-distribution for the student RPN. In contrast, the attention maps from self-generated responses, even if imperfect, are inherently aligned with its own internal visual grounding mechanisms. This creates a more consistent and attainable distillation target, especially in our data-efficient setting. This self-sufficient framework thus removes any dependency on external models or annotated data.